OpenStudy (anonymous):

how to find the equation of a tangent line y= sqrt 1+4 sinx at point (0,1)

OpenStudy (anonymous):

The tangent line \(y\) of a function \(f(x)\) at the point \((a,f(a))\) is: \[ y-f(a)=f'(a)(x-a) \]

OpenStudy (anonymous):

In this case \(a=0,f(a)=1\). You need to find the derivative of \(f(x)=\sqrt{1+4\sin(x)}\)

OpenStudy (anonymous):

okay, so far i have y'= 1/2 (1+4sinx)^-1/2 (cosx) but idk if the cosx is right or what to do then.

OpenStudy (anonymous):

So you want help finding the derivative?

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

I would do implicit differentiation: \[ \begin{array}{} y&=&\sqrt{1+4\sin(x)}\\ y^2&=&1+4\sin(x)\\ y^2-1&=&4\sin(x)\\ 2yy'&=&4\cos(x)\\ y' &=&\frac{4\cos(x)}{2y}\\ y' &=&\frac{4\cos(x)}{2\sqrt{1+4\sin(x)}} \end{array} \]

OpenStudy (anonymous):

You can also use chain rule: \[ y=(f\circ g)\\ f=\sqrt{x}\\ g=1+4\sin(x)\\ \]\[ f'=\frac{1}{2\sqrt{x}}\\ g'=4\cos(x)\\ y'=f(g)\cdot g'=\frac{4\cos(x)}{2\sqrt{1+4\sin(x)}} \]

OpenStudy (anonymous):

@ansleyb Do you think you can finish the rest?

OpenStudy (anonymous):

yah thank you